### v2

# Geometry induced net spin polarization of $d$-wave altermagnets

Abstract
Altermagnets exhibit spin-split electronic bandstructures despite having zero net magnetization, making them attractive for field-free spintronic applications. In this work, we show that a finite rectangular altermagnetic sample can acquire a net spin polarization purely due to its geometry. This effect arises from the interplay between the anisotropic, spin-resolved Fermi contours of an altermagnet and the discrete sampling of momentum space imposed by unequal sample dimensions. By explicitly counting occupied states, we demonstrate that rectangular samples with $L_{x}\neq L_{y}$ host a finite spin polarization, which vanishes in the symmetric limit $L_{x}=L_{y}$ and in the thermodynamic limit. We further show that this geometry-induced spin polarization can be directly probed in transport measurements. In the tunneling regime, the charge and the spin conductances exhibit characteristic patterns as a function of sample dimensions, faithfully reflecting the underlying spin polarization. In addition, transport across ferromagnet–altermagnet–ferromagnet junctions reveals an asymmetric magnetoresistance with respect to reversal of the Zeeman field, providing an independent transport signature of the finite spin polarization. Our results establish geometry as an effective control parameter for spin polarization in altermagnets and suggest a viable route for exploiting finite-size effects in mesoscopic altermagnetic spintronic devices.
I Introduction
Altermagnets (AMs) are a recently identified class of magnetic materials that exhibit spin-split electronic bandstructures despite having zero net magnetization, originating from magnetic order that breaks time-reversal symmetry while preserving combined symmetries [10]. In particular, $d$-wave altermagnets host a $d$-wave magnetic order, drawing a close analogy to unconventional superconductors with $d$-wave pairing symmetry. The absence of net magnetization is a defining characteristic of AMs and renders them especially attractive for field-free spintronic applications, where stray magnetic fields are undesirable [1, 13, 9, 8].
The spin-split nature of the electronic structure in AMs has led to a growing interest in their transport properties and device potential. Notably, it enables the generation of spin currents in hybrid junctions involving normal metals, even though both constituent materials individually possess zero net spin polarization [2]. Furthermore, the spin-dependent band anisotropy gives rise to orientation-dependent Andreev reflection at AM–superconductor interfaces [6, 11] and can strongly enhance crossed Andreev reflection in altermagnet-based heterostructures by suitably controlling the relative orientations of the magnetic order [3]. Electron transport across AMs has been explored in a variety of settings, both theoretically [2, 12] and experimentally [7, 5], establishing AMs as a promising platform for spin-dependent transport phenomena.
Beyond these effects, the anisotropic nature of the spin-resolved bandstructures in altermagnets provides an additional and largely unexplored route to controlling spin polarization. In particular, when the electronic spectrum is sampled discretely, as in finite-sized systems, the anisotropy of the Fermi contours can lead to an unequal population of the two spin species. This opens up the possibility of inducing a net spin polarization purely through sample geometry, without invoking external magnetic fields or intrinsic magnetization. Related finite-size effects have been identified in altermagnetic strips, where a net local magnetization develops near the edges [4]. In contrast, we focus on a geometry-induced spin polarization that arises from the global shape and aspect ratio of the sample and manifests itself in transport measurements. Specifically, we analyze this effect in finite altermagnetic samples and propose experimentally accessible schemes to detect it, highlighting finite-size and shape effects as effective control parameters in altermagnet-based spintronic devices.
II Spin polarization of a two-dimensional rectangular AM
In this section, we demonstrate that a two-dimensional altermagnet (AM) with a rectangular geometry can acquire a finite net spin polarization. The low-energy Hamiltonian describing the AM is given by
$H$ $=$ $-ta^{2}(\partial_{x}^{2}+\partial_{y}^{2})\sigma_{0}+t_{J}a^{2}(\partial_{x}^{2}-\partial_{y}^{2})\sigma_{z},$ (1)
where $t>0$ sets the kinetic energy scale, $t_{J}$ ($0<t_{J}<t$) characterizes the strength of the altermagnetic order [10] which results in spin-split, anisotropic bandstructure, $a$ is a fundamental length scale corresponding to the lattice spacing, and $\sigma_{i}$ are Pauli matrices in spin space.
We consider a rectangular two-dimensional AM sample of length $L_{x}$ and width $L_{y}$. Imposing periodic boundary conditions along both directions, the allowed wavevectors are $k_{x}=n_{x}\pi/L_{x}$ and $k_{y}=n_{y}\pi/L_{y}$, where $n_{x},n_{y}$ are integers. The dispersion relations for electrons with spin $\sigma$ ($\sigma=\uparrow,\downarrow$) are then given by
$E_{\sigma}(\vec{k})$ $=$ $ta^{2}(k_{x}^{2}+k_{y}^{2})-s_{\sigma}t_{J}a^{2}(k_{x}^{2}-k_{y}^{2}),$~~ (2)
with $s_{\uparrow}=+1$ and $s_{\downarrow}=-1$.
For a given Fermi energy $E_{F}$, we numerically evaluate the total number of occupied states $N_{\sigma}$ for each spin species by counting the allowed $(k_{x},k_{y})$ modes satisfying $E_{\sigma}(k_{x},k_{y})<E_{F}$. Figure 1 shows the normalized spin polarization $(N_{\uparrow}-N_{\downarrow})/(N_{\uparrow}+N_{\downarrow})$ as a function of the sample dimensions $(L_{x},L_{y})$ for $t_{J}=0.8t$ and $E_{F}=0.6t$.
Figure 1: Normalized spin polarization $(N_{\uparrow}-N_{\downarrow})/(N_{\uparrow}+N_{\downarrow})$ versus $(L_{x},L_{y})$ for $t_{J}=0.8t$ and $E_{F}=0.6t$ with periodic boundary conditions. The samples with $L_{x}\neq L_{y}$ exhibit nonzero spin polarization.
For $L_{x}=L_{y}$, the net spin polarization vanishes. When $L_{x}\neq L_{y}$, the geometric symmetry between the two directions is broken and one direction is preferentially selected. In this case, the anisotropic bandstructure of one spin species, which is elongated along the $x$ direction, accommodates a larger number of allowed $k_{x}$ states for a fixed $k_{y}$ satisfying $E(\vec{k})<E_{F}$, compared to the bandstructure of the opposite spin, which is elongated along the $y$ direction. As a result, an imbalance in the occupation of the two spin species emerges, leading to a finite net spin polarization.
For a rectangular sample with $L_{x}\neq L_{y}$, the discretization of momentum space is anisotropic, with spacings $\Delta k_{x}\sim\pi/L_{x}$ and $\Delta k_{y}\sim\pi/L_{y}$. Because the spin-resolved Fermi contours in an altermagnet are anisotropic and elongated along orthogonal directions for opposite spin species, this geometry-induced discretization affects the two spins differently. For the spin whose Fermi contour is elongated along the $x$ direction, a larger $L_{x}$ implies a finer sampling of momenta along $k_{x}$, allowing more states to fall within the Fermi contour for a given range of $k_{y}$. In contrast, the opposite spin, whose Fermi contour is elongated along the $y$ direction, benefits less from this enhanced resolution in $k_{x}$.
As a result, the imbalance in the number of occupied states depends sensitively on the aspect ratio of the sample, and the normalized spin polarization increases with $|L_{x}/L_{y}-1|$, vanishing in the symmetric limit $L_{x}=L_{y}$. However, it is important to note that the spin species with a Fermi contour elongated along the $y$ direction samples a correspondingly larger range of $k_{y}$, which compensates for the reduced resolution along $k_{x}$. Consequently, when the system size becomes large, the total number of states for the two spin species scales similarly, and the net spin polarization per unit area approaches zero. This highlights that the geometry-induced spin polarization is predominantly a finite-size effect, most pronounced in small samples.
When periodic boundary conditions are imposed, the allowed momenta are quantized as $k_{j}=n_{j}\pi/L_{j}$ for $j=x,y$, where $n_{j}$ takes integer values, including zero. The condition $E(\vec{k})<E_{F}$ then translates into $|n_{j}|<k_{F,j}L_{j}/\pi$ for modes with $n_{j^{\prime}}=0$, where $j^{\prime}=y,x$ for $j=x,y$, and the spin-dependent Fermi wavevectors are given by $k_{F,x}=\sqrt{E_{F}/[(t-s_{\sigma}t_{J})a^{2}]}$ and $k_{F,y}=\sqrt{E_{F}/[(t+s_{\sigma}t_{J})a^{2}]}$. Consequently, the square-like features observed in Fig. 1 originate from modes with either $n_{x}=0$ or $n_{y}=0$. This behavior changes qualitatively when open boundary conditions are imposed, for which $n_{j}$ is restricted to positive integers and the $n_{j}=0$ modes are excluded. As a result, the square-like pattern disappears in the normalized spin polarization plotted as a function of sample dimensions, as shown in Fig. 2. This distinction is particularly relevant for experiments, since periodic boundary conditions are an idealization and realistic finite samples are governed by open boundary conditions, for which the square-like features are absent.
Figure 2: (a) Normalized spin polarization $(N_{\uparrow}-N_{\downarrow})/(N_{\uparrow}+N_{\downarrow})$ (b) $(N_{\uparrow}+N_{\downarrow})$ (c) $(N_{\uparrow}-N_{\downarrow})$, versus the sample dimensions $(L_{x},L_{y})$. The samples with $L_{x}\neq L_{y}$ exhibit nonzero spin polarization.
III Transport probe of geometry-induced spin polarization
An experimental probe of the geometry-induced spin polarization in AMs is essential not only for its direct detection but also for assessing its potential utility in spintronic applications. To this end, we propose a transport-based setup in which a rectangular AM sample is sandwiched between two normal-metal electrodes with rectangular cross sections, as shown in Fig. 3. We demonstrate that electron transport through the AM in this geometry exhibits characteristic signatures that directly reflect the presence of a net spin polarization in the AM.
Figure 3: A thin rectangular slab of AM connected to NM electrodes with rectangular cross-section of same size. The dimensions of the AM slab is $L_{x}\times L_{y}\times d$. The NM leads are extended along $z$-direction and have same cross sectional dimensions as that of AM.
The Hamiltonian to study transport in such a setup is given by
$H$ $=$ $\begin{cases}-t_{n}a^{2}(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})\sigma_{0},~~~{\rm for~~~}|z|>d/2,\\ ~~\\ -ta^{2}(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})\sigma_{0}+t_{J}a^{2}(\partial_{x}^{2}-\partial_{y}^{2})\sigma_{z}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~{\rm for~~}|z|<d/2,\end{cases}$ (3)
where $d$ is the thickness of the AM slab and $t_{n}$ is the kinetic energy scale in NM electrodes. The boundary conditions to solve the scattering problem at $z=\pm d/2$ are given by
$\partial_{z}\psi|_{z_{0}^{+}}-\partial_{z}\psi|_{z_{0}^{-}}$ $=$ $q_{0}\psi(z_{0}),~~~~{\rm for}~~z_{0}=\pm d/2,$ (4)
in addition to the continuity of $\psi$. For all $z$, the system is confined to $0<x<L_{x}$ and $0<y<L_{y}$. We impose open boundary conditions along the $x$ and $y$ directions, as these provide a more realistic description of a finite sample. Along the $z$ direction, the AM slab is subject to open boundary conditions. As a consequence, the lowest allowed transverse mode acquires a finite energy shift $ta^{2}\pi^{2}/d^{2}$ compared to the strictly two-dimensional case, where $d$ denotes the thickness of the AM slab.
Figure 4: (a) Charge conductance in units of $e^{2}/h$ and (b) spin conductance in units of $e/2$ across AM versus the system size $(L_{x},L_{y})$ for $t_{n}=0.2t$, $t_{J}=0.8t$, and $d=5a$ for the system with open boundary conditions along $x$ and $y$.
The transport properties are evaluated using the Landauer formalism, with the conductance calculated independently in the two spin sectors. The sum of the two contributions defines the charge conductance $G$, while their difference yields the spin conductance $G^{s}$. In Fig. 4, we plot (a) $G$ and (b) $G_{s}$ versus $(L_{x},L_{y})$ for $q_{0}=10$, $E_{F}=t$, with $t_{n}=0.2t$, $t_{J}=0.8t$, and $d=5a$. The hopping parameter $t_{n}$ in the normal-metal leads is chosen such that their Fermi surface fully encloses that of the AM, ensuring that all relevant AM bands participate in transport.
As evident from Fig. 4, both the charge conductance $G$ and the spin conductance $G_{s}$ exhibit extrema along contours that coincide with the transitions in the number of occupied states shown in Fig. 2. Along these contours, the energy of a state in the isolated AM sample matches the energy of electrons incident from the leads, giving rise to transmission resonances. We emphasize that this direct correspondence between spin polarization and transport signatures emerges only in the tunneling limit.
Next, we study transport across an FM–AM–FM junction, where the normal-metal electrodes in Fig. 3 are replaced by ferromagnetic (FM) leads. The Hamiltonian for the FM electrodes is
$H=-t_{n}a^{2}(\partial_{x}^{2}+\partial_{y}^{2}+\partial_{z}^{2})\sigma_{0}-b\,\sigma_{z},$
valid in the regions $|z|>d/2$. Here, $b$ denotes the Zeeman energy scale, proportional to the magnetization of the FM leads. All other details of the setup remain unchanged.
Figure 5: Magnetoresistance (MR) of the FM–AM–FM junction. (a) MR as a function of the Zeeman energy $b$ for different system sizes. Each curve corresponds to a distinct choice of $(L_{x},L_{y})/a$, as indicated in the legend. (b) MR versus $b$ for systems with the same aspect ratio, $(L_{x},L_{y})=n(4,10)a$, for different values of $n$ shown in the legend. (c) The asymmetry measure $\mathrm{MR}(b)-\mathrm{MR}(-b)\,]/2b$ as a function of the characteristic system size $L_{0}$, for $b=0.1t$. It can be seen that in the limit of large system sizes the asymmetry measure tends to zero. All other parameters are the same as in Fig. 4.
We compute the conductance $G(b)$ as a function of the Zeeman energy $b$ and characterize the transport response via the magnetoresistance (MR), defined as $\mathrm{MR}=\delta R(b)/R(0)$, where $\delta R(b)=R(b)-R(0)$ and $R(b)=1/G(b)$ is the resistance at Zeeman energy $b$. In Fig. 5(a), we plot the MR as a function of $b$ for several choices of the aspect ratio $L_{x}/L_{y}$, while keeping the cross-sectional area fixed. For rectangular samples with $L_{x}\neq L_{y}$, the MR is clearly asymmetric under $b\to-b$, reflecting the presence of a nonzero net spin polarization in the altermagnet. In contrast, for $L_{x}=L_{y}$, the MR is symmetric in $b$, consistent with the absence of net spin polarization.
A symmetric MR–$b$ curve indicates the absence of net spin polarization, whereas an asymmetric dependence signals a finite spin polarization. In Fig. 4(b), we plot the MR as a function of $b$ for several system sizes sharing the same aspect ratio. As the system size increases, the slope of the MR curve at $b=0$ is progressively reduced. We quantify this asymmetry using the slope at $b=0$, which may be approximated by $\mathrm{MR}(b)-\mathrm{MR}(-b)\,]/2b$ for small $b$. Figure 4(c) shows this asymmetry measure as a function of system size for a fixed aspect ratio. In the limit of large system sizes, the asymmetry tends to zero, indicating that the net spin polarization in altermagnets is predominantly a finite-size effect.
IV Summary and conclusions
In this work, we have shown that a finite rectangular altermagnetic sample can exhibit a net spin polarization purely due to its geometry. This effect originates from the anisotropic, spin-split bandstructure of altermagnets, combined with the discrete sampling of momentum space imposed by finite sample dimensions. For rectangular samples with unequal side lengths, the anisotropic discretization of $\vec{k}$ space affects the two spin species differently, leading to an imbalance in the number of occupied states and hence a finite spin polarization.
We demonstrated that this geometry-induced spin polarization can be directly detected in transport measurements. In the tunneling regime, the spin polarization inferred from state counting is faithfully reflected in the charge and spin conductances which exhibit characteristic patterns as a function of the sample dimensions. These features establish a direct correspondence between the underlying band-structure anisotropy and experimentally accessible transport observables.
Furthermore, we investigated transport across FM–AM–FM junctions and showed that the presence of a net spin polarization in the AM manifests itself as an asymmetry in the magnetoresistance under reversal of the Zeeman field in the ferromagnetic leads. This asymmetry disappears for square samples, consistent with the vanishing of the geometry-induced spin polarization when $L_{x}=L_{y}$. We also found that the asymmetry decreases with increasing system size, reflecting the fact that the spin polarization per unit area vanishes in the thermodynamic limit.
Our results highlight a route to generating and detecting spin polarization in altermagnets without invoking external magnetic fields or intrinsic magnetization, relying instead on sample geometry and band-structure anisotropy. This finite-size effect should be most relevant for mesoscopic devices and may offer a useful design principle for altermagnet-based spintronic applications.
Acknowledgements.
The author thanks Jacob Linder and Pavlo Sukhachov for useful discussions. The author thanks Science and Engineering Research Board (now Anusandhan National Research Foundation) Core Research grant (CRG/2022/004311) and University of Hyderabad for financial support.